Sin^2(x+30)/3 + cos^2(x/2)= tan 45, find tanX value

Grade:10

1 Answers

Aditya Gupta
5 years ago
we know that tan45=1
so Sin^2(x+30)/3 = 1 – cos^2(x/2)= sin^2(x/2)
so so x/3+10= npi+-x/2
so now tanx can be found out as we have found the solution for x